# Magnetic field amplification and decay in cosmic string wakes

### v1

@@ 1 Introduction
In the literature, the magnetic field in the wake region is generated due to the presence of density inhomogeneities in the plasma. These density inhomogeneities come from the clustering of different particles around the cosmic string. The clustering of these particles mostly depend on their masses abhisek . A long and narrow wake is formed. In our numerical simulations, we have found that when the velocity of the cosmic string is low, the wake structure is diffuse. A well developed wake structure is obtained for higher Mach numbers. For simplicity, we consider an ideal fluid approximation without any dissipative forces. The generated vorticity in the plasma is thus conserved. We proceed to study the evolution of the field over time in the absence of any turbulent forces.
@@
In section II, we discuss the spacetime around a cosmic string and discuss how vorticity is generated in a cosmic string wake. In section III, we discuss the amplification of the magnetic field due to the deformation of the magnetic flux lines. In Section IV, we show the decay of the magnetic field in presence of a sheared magnetic field. We also show that in the presence of a sheared magnetic field, the amplification or decay depends on the orientation and the lengthscale of the perturbative field. Section V we summarize the work and discuss why it is important to understand the evolution of magnetic fields in the cosmic string wakes for current observational signatures.
@@ 2 Vorticity in a cosmic string wake
The only difference comes as the particles move past the string. The conical nature of the metric gives a velocity perturbation to the particle towards the cosmic string. This velocity perturbation is proportional to the deficit angle $\Delta\theta$ of the string. It is given by $\delta v\approx\delta\theta v_{s}\gamma_{s}$, where, $v_{s}$ is the cosmic string velocity and $\gamma_{s}$ is the Lorentz gamma factor corresponding to the moving cosmic string. This causes the particle trajectories to bend towards the moving cosmic string. Though small, $\delta\theta$ is always non-zero, hence, ×$\triangledown\times\mathbf{v}$ is non-zero. A rotational flow is thus generated. The local vorticity of the plasma is defined by,
$\boldsymbol{\omega_{L}}=\triangledown\times\mathbf{v}$ (5)
Here, $\mathbf{v}$ is the particle velocity. Since the deficit angle is small, the vorticity is also generated at smaller lengthscales.
In the case of magnetohydrodynamics, the vorticity may be enhanced by the presence of the charged paricles in the plasma. For the ideal fluid approximation, the magnetic field lines are frozen into the flowing fluid. It is possible to define a generalized vorticity for the fluid which considers the magnetic field. Charged particles move in rotational motion in a magnetic field, hence the vorticity is enhanced. The generalized vorticity is defined by mahajan ,
$\boldsymbol{\omega_{G}}=\triangledown\times\mathbf{v}+\frac{q\mathbf{B}}{c}$ (6)
For an ideal fluid, the vorticity generated in the fluid will always be conserved. In most systems with ideal fluids, it is the boundary conditions which lead to the generation of vorticity but in the case of the particles moving close to a cosmic string, an angular momentum is generated as the particles orbit the string. The orbital angular momentum then leads to a non-zero ×$\triangledown\times\mathbf{v}$ in the fluid. So, the vorticity is generated in the wake of a cosmic string as it moves through the cosmic fluid. In the case of hydrodynamics, this local vorticity is small but in the case of magnetohydrodynamics, this vorticity is important for identifying lengthscales of the magnetic field perturbations. We show later on why “coarse-graining” the fields and the smallest lengthscale are important for the growth or decay of the magnetic field.
@@ 2.1 Numerical Simulation of the vorticity
We use the publicly available OpenMHD code openmhd to simulate the vorticity in the cosmic string wakes. We use the basic code and use open boundary conditions in the direction of fluid flow. In the basic code, the MHD equations along with the equation of state, relating the hydrodynamic pressure and the density, $P_{H}=f(\rho)$ are solved using a perturbative method. For the case of magnetogydrodynamics, the total pressure ($P$) consists of the hydrodynamic pressure and the pressure due to the magnetic field mkverma . The fields ($\mathbf{v}$, $\mathbf{B}$, $P$, $\rho$) mentioned previously in the MHD equations can be split into a background field and a perturbative field as, $\mathbf{v}=\mathbf{v_{0}}+\mathbf{v_{1}}$, $\mathbf{B}=\mathbf{B_{0}}+\mathbf{B_{1}}$, $P=P_{0}+P_{1}$ and $\rho=\rho_{0}+\rho_{1}$. Here $(\mathbf{v_{0}}$, $\mathbf{B_{0}}$, $P_{0}$, $\rho_{0})$ are the background fields, and $(\mathbf{v_{1}}$, $\mathbf{B_{1}}$, $P_{1}$, $\rho_{1})$ are the perturbative fields. A Galilean transformation is used to shift to the background flow frame, thus, $\mathbf{v_{0}}=0$, $\mathbf{v_{1}}=\mathbf{v}$ which means that the velocity field $\mathbf{v}$ is the perturbative field. These are then substituted in the MHD equations mentioned previously, and only the first order terms are retained, assuming the perturbations to be small. The linearized equations are given by,
@@
The basic code is available in one and two dimensions. We have used the 2D code and modified it by giving a velocity perturbation to the particles as they cross the cosmic string. The fluid flow is taken to be along the $x$-axis. In the $y$-direction, we have used fixed boundary conditions. The cosmic string is aligned with the $z$-axis, it is identified in the flow by the position at which an initial velocity perturbation is given to the plasma particles. All the particles are subject to this velocity perturbation which will depend on the deficit angle of the cosmic string metric. Thus, at the initial time, plasma particles crossing a certain position ($x$ coordinate) are all given a velocity perturbation both in the $x$-and $y$-direction; $v_{x}=-v_{0}Cos\theta$ and $v_{y}=-v_{0}Sin\theta$. All other particles move with velocity $-v_{0}$. This is in the reference frame of the cosmic string. As the particles flow, after some time the vortex structure is formed. Fig.1 shows a streamline plot of the vortex structure generated in a cosmic string wake.
@@
For a better understanding of the simulation, a larger deficit angle is taken so that the vorticity is clearly visible. However, it is possible to scale the results to the cosmic string lengthscales. As mentioned before, the deficit angle for an actual cosmic string is very small, in dimensionless parameters the perturbation would be of the order of the deficit angle with a value of $10^{-5}$ gangui . Based on this, the lengthscale of the vortex will be of the order of $10^{-5}Mpc$ or smaller. The net vorticity over a certain region is calculated and in the absence of a magnetic field, we find that the vorticity is conserved throughout the flow. This is an important check, as we have mentioned that we do not consider any dissipative forces in the system, so the energy conservation will also mean the conservation of vorticity in the flow.
@@ 3 Sheared magnetic fields and field amplification
The velocity perturbation on the particles is along the $y$-direction, so the particles move towards the conical space behind the cosmic string. We have given a sheared magnetic field as the magnetic field perturbation. Though we are perturbing the $x$ component, the variation is along the $y$-direction. As the magnetic field lines are being pushed in the $y$-direction by the velocity flow, so our perturbation is enhancing the effect of the compression of the magnetic field lines due to the fluid flow. We find that we get an amplification of the field for all values of $\alpha_{1}$.

### v2

@@ 1 Introduction
In the literature, the magnetic field in the wake region is generated due to the presence of density inhomogeneities in the plasma. These density inhomogeneities come from the clustering of different particles around the cosmic string. The clustering of these particles mostly depend on their masses abhisek . A long and narrow wake is formed. In our numerical simulations, we have found that when the velocity of the cosmic string is low, the wake structure is diffuse. A well developed wake structure is obtained for higher Mach numbers. For simplicity, we consider an ideal fluid approximation without any dissipative forces. To generate some vorticity in the wake, we give a small velocity perturbation to the particle velocities in both the directions in the plane of the wake. The small vorticity generated in the plasma is conserved. Since there are no turbulent forces, we do not study the evolution of the vorticity with time. Our main aim is to study the evolution of the magnetic field over time in the absence of any turbulent forces.
@@
In section II, we discuss the spacetime around a cosmic string and discuss how we generate vorticity in the cosmic string wake structure. In section III, we discuss the amplification of the magnetic field due to the deformation of the magnetic flux lines. In Section IV, we show the decay of the magnetic field in presence of a sheared magnetic field. We also show that in the presence of a sheared magnetic field, the amplification or decay depends on the orientation and the lengthscale of the perturbative field. Section V we summarize the work and discuss why it is important to understand the evolution of magnetic fields in the cosmic string wakes for current observational signatures.
@@ 2 Vorticity in a cosmic string wake
The only difference comes as the particles move past the string. The conical nature of the metric gives a velocity perturbation to the particle towards the cosmic string. This velocity perturbation is proportional to the deficit angle $\Delta\theta$ of the string. It is given by $\delta v\approx\delta\theta v_{s}\gamma_{s}$, where, $v_{s}$ is the cosmic string velocity and $\gamma_{s}$ is the Lorentz gamma factor corresponding to the moving cosmic string. But this perturbation does not generate a local vorticity. The cosmic string metric is conical, which means that it is locally flat. From the reference frame of the particle, the path of the particle is always a straight line unless it is acted upon by some external force. Since the particles always move in a straight line, no local vorticity can be generated by the particles.
From an external observer’s reference frame however, the trajectories of the particles are bent. This gives rise to the overdensity of the fluid behind a moving cosmic string. However, massive particles which constitute the plasma of the early universe do have an angular momentum if the cosmic strings are Abelian Higgs strings hartmann . It has been shown that massive test particles can have a bound orbit or an unbound circular orbit around a Abelian Higgs cosmic string. It is therefore possible that the trajectory of the particle may not be a straight line in the reference frame of the particle. In this work, we do not specify the nature of the cosmic string, but we assume that the velocity of the particle moving in the plasma is affected by the collisions with other plasma particles. We consider a hydrodynamic approach where the velocity space of the number density of the particles in a plasma is treated as a distribution function. We generate a local vorticity in the wake behind the string by giving a perturbation to the velocity in both the $x$ and $y$ directions. These are the two directions in which the plane of the wake is considered. We consider these perturbations to the particle velocity to be of the order of the $\delta v$.
The perturbation to the velocity in both the directions may be due to various reasons. There are density inhomogeneities generated at several lengthscales in the wake of the cosmic string. The different species of particles in the wake cluster at different lengthscales. The collision between these particles may generate velocity perturbations in the plasma. Particle clustering non-uniformly close to an Abelian Higgs string has been discussed in previous hydrodynamic studies of wake structures sornborger , so small perturbations to the velocity of the particles is quite possible. This is why we give a small initial perturbation to the velocity in both the planar directions of the wake. This leads to the generation of a local vorticity in the plasma. This is shown if fig 1.
Vorticity in magnetohydrodynamic wakes are more common than in hydrodynamic wakes. This is because of the presence of the magnetic field. If we consider the trajectory of a charged particle, such as an electron in the magnetized plasma in a flat space time, the equation of motion is given by,
$m_{e}[\frac{\partial\mathbf{v_{e}}}{\partial t}+(\mathbf{v_{e}}.\nabla)\mathbf{v_{e}}]=\frac{-\nabla P_{e}}{n_{e}}+e_{e}\mathbf{E}+\frac{q}{c}\mathbf{v_{e}}\times\mathbf{B}$ (5)
Here the subscript ${e}$ denotes the electron as the charged particle but such an equation can be written down for any species of charged particles. Here $m_{e}$ denotes the mass of the particle and $P_{e}$, the hydrodynamic pressure. $n_{e}$ is the number density of the charged particles. In the presence of a magnetic field therefore, it is possible to define a generalized vorticity for the fluid. Charged particles move in rotational motion in a magnetic field and the generalized vorticity is defined by mahajan ,
$\boldsymbol{\omega_{G}}=\triangledown\times\mathbf{v_{e}}+\frac{q\mathbf{B}}{c}$ (6)
This equation is valid as long as the density is a function of pressure. In the case of the ion being the charged particle, similar equations can be written down. To obtain the one fluid MHD equations, some approximations have to be made. The mass of the electron is considered to be negligible compared to the mass of the ions. The plasma is assumed to be quasi neutral, $|n_{e}-n_{i}|<<n_{e}$. It is also assumed that the dispacement current from the Maxwell’s equation is negligible. Based on these assumptions, the ideal one fluid MHD equations can be used to study the evolution of the magnetic field.
As is well known, for an ideal fluid, the vorticity generated in the fluid will always be conserved. So once we generate the vorticity due to the velocity perturbations in the plasma it remains constant. After generating the vorticity, we use the one fluid model of the MHD equations to evolve the different fields (pressure, magnetic field etc). The vorticity remains constant throughout the evolution of the wake structure and therefore we do not study it’s evolution with time. Moreover, since the magnetic field lines are frozen in the plasma, the evolution of the equations of motion of the charged particles will be related to both the magnetic pressure as well as the hydrodynamic pressure in the fluid. We show later on why “coarse-graining” the fields and the smallest lengthscale are important for the growth or decay of the magnetic field in the magnetized plasma.
@@ 2.1 Numerical Simulation of the vorticity
We use the publicly available OpenMHD code openmhd for our simulations. We use the basic code and use open boundary conditions in the direction of the fluid flow. In the basic code, the MHD equations along with the equation of state, relating the hydrodynamic pressure and the density, $P_{H}=f(\rho)$ are solved using a perturbative method. For the case of magnetogydrodynamics, the total pressure ($P$) consists of the hydrodynamic pressure and the pressure due to the magnetic field mkverma . The fields ($\mathbf{v}$, $\mathbf{B}$, $P$, $\rho$) mentioned previously in the MHD equations can be split into a background field and a perturbative field as, $\mathbf{v}=\mathbf{v_{0}}+\mathbf{v_{1}}$, $\mathbf{B}=\mathbf{B_{0}}+\mathbf{B_{1}}$, $P=P_{0}+P_{1}$ and $\rho=\rho_{0}+\rho_{1}$. Here $(\mathbf{v_{0}}$, $\mathbf{B_{0}}$, $P_{0}$, $\rho_{0})$ are the background fields, and $(\mathbf{v_{1}}$, $\mathbf{B_{1}}$, $P_{1}$, $\rho_{1})$ are the perturbative fields. A Galilean transformation is used to shift to the background flow frame, thus, $\mathbf{v_{0}}=0$, $\mathbf{v_{1}}=\mathbf{v}$ which means that the velocity field $\mathbf{v}$ is the perturbative field. These are then substituted in the MHD equations mentioned previously, and only the first order terms are retained, assuming the perturbations to be small. The linearized equations are given by,
@@
The basic code is available in one and two dimensions. We have used the 2D code and modified it by giving a velocity perturbation in both the directions of the plane of the wake to the particles as they cross the cosmic string. The fluid flow is taken to be along the $x$-axis. In the $y$-direction, we have used fixed boundary conditions. The cosmic string is aligned with the $z$-axis, it is identified in the flow by the position at which an initial velocity perturbation is given to the plasma particles. Thus, at the initial time, plasma particles crossing a certain position ($x$ coordinate) are all given a velocity perturbation both in the $x$-and $y$-direction; $v_{x}=-v_{0}Cos\theta$ and $v_{y}=-v_{0}Sin\theta$. All other particles move with velocity $-v_{0}$. This is in the reference frame of the cosmic string. As the particles flow, after some time, the vortex structure is formed. Fig.1 shows a streamline plot of the vortex structure generated in a cosmic string wake.
@@
For a better understanding of the simulation, a larger deficit angle is taken so that the vorticity is clearly visible. A larger deficit angle will result in a bigger dimension of the wake. We prefer to keep the order of the velcocity perturbations that we give to the particles of the same order as the velocity perturbation due to the cosmic string metric. As mentioned before, the deficit angle for an actual cosmic string is very small, in dimensionless parameters the perturbation would be of the order of the deficit angle with a value of $10^{-5}$ gangui . If we give a perturbation of the same order to the velocity along both the dimensions of the wake, the lengthscale of the vortex will be of the order of $10^{-5}Mpc$ or smaller. In our simulations, the net vorticity over a certain region is calculated and in the absence of a magnetic field, we find that the vorticity is conserved throughout the flow. This is an important check, as we have mentioned that we do not consider any dissipative forces in the system, so the energy conservation will also mean the conservation of vorticity in the flow.
@@ 3 Sheared magnetic fields and field amplification
There is a velocity perturbation on the particles along the $y$-direction, so the particles move towards the conical space behind the cosmic string. We have given a sheared magnetic field as the magnetic field perturbation. Though we are perturbing the $x$ component, the variation is along the $y$-direction. As the magnetic field lines are being pushed in the $y$-direction by the velocity flow, so our perturbation is enhancing the effect of the compression of the magnetic field lines due to the fluid flow. We find that we get an amplification of the field for all values of $\alpha_{1}$.